The population of fish in a particular pond is known to have a mean length µ = 15 cm with standard deviation σ = 5. You catch fifty fish from the pond and measure their lengths. What is the probability that the average length of those fish (i.e. sample mean X ) is between 14 cm and 16.5 cm?


1
Expert's answer
2021-11-30T10:32:47-0500

Let "X=" sample mean: "X\sim N(\mu, \sigma^2/n)."

Given "\mu=15, \sigma=5, n=50."


"P(14<X<16.5)=P(X<16.5)-P(X\leq 14)"

"=P(Z<\dfrac{16.5-15}{5/\sqrt{50}})-P(X\leq \dfrac{14-15}{5/\sqrt{50}})"

"\approx P(Z<2.1213)-P(X\leq -1.4142)"

"\approx0.98305-0.07865\approx 0.9044"

The probability that the average length of those fish (i.e. sample mean X ) is between 14 cm and 16.5 cm is 0.9044.